1. Solve the following pair of simultaneous equations.
6f + e =23
2e - f = 7

Solve question 1. above graphically, i.e. using straight line graphs. Show all relevant points clearly.

Answer by mananth(16946)

6 f + 1 e = 23 .............1
-1 f 2 e = 7 .............2
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when f= 0 e= 23 ,( 0 , 23 )
when e = 0 f= 23 ,( 3.83 , 0 )
Equation 2
when f= 0 e= 3.5 ( 0 , 3.5 )
when e = 0 f= -7 ( 0 , -7 )



You will observe that the point of intersection is 3 , 5
The solution of the system is ( 3 , 5 )